# Follow Everything: A Leader-Following and Obstacle Avoidance Framework with Goal-Aware Adaptation

### v1

@@ III-B Goal States Adaptation and Trajectory Planning
Unlike conventional approaches that select the local goal as the intersection point between the global path and the costmap boundary, the chasing state extends this local point along the costmap edge to form a goal line of length $L_{t}$, which is proportional to the distance between the robot $p_{r}$ and the leader $\bar{p}_{l}$:
$L_{t}=\alpha(|p_{r}-\bar{p}_{l}|-\frac{W_{map}}{2})$ (6)
@@
The whole goal line is separated at the locations where it intersects with obstacles, resulting in several sub-goal line sets $\mathcal{G}_{i}$, which are then added as additional nodes to the topological graph along with their corresponding connections.
Several generalized trajectories can be generated over the topological graph using depth-first search, with each one $\mathcal{T}$ following the form:
@@
Since each obstacle node $\mathcal{N}_{i}$ can be detoured in either a clockwise or counter-clockwise direction, a generalized trajectory $\mathcal{T}$ containing $n$ nodes can yield up to $2^{n}$ distinct trajectory candidates $\tau$. To eliminate semantically redundant trajectories that represent the same detour behavior, we apply the concept of trajectory homotopy classes [27, 28] to prune equivalent candidates. The remaining trajectories are then optimized in parallel.
Following the time-optimal optimization framework [29], in addition to standard kinematic, dynamic (with the max speed $V^{max}$), and obstacle-avoidance constraints, an additional goal constraint is introduced here:
$\begin{gathered}|\overrightarrow{\mathcal{G}_{k}^{start}\;\mathcal{G}_{k}^{end}}\times\overrightarrow{\mathcal{G}_{k}^{start}\;p_{\tau}^{end}}|=0\\ 0\leq\overrightarrow{\mathcal{G}_{k}^{start}\;\mathcal{G}_{k}^{end}}\cdot\overrightarrow{\mathcal{G}_{k}^{start}\;p_{\tau}^{end}}\leq|\overrightarrow{\mathcal{G}_{k}^{start}\;p_{\tau}^{end}}|\end{gathered}$ (8)
where $\mathcal{G}_{k}^{start}$ and $\mathcal{G}_{k}^{end}$ denote the endpoints of a goal set $\mathcal{G}_{k}$. The trajectory end point $p_{\tau}^{end}$ is allowed to slide along $\mathcal{G}_{k}$, enhancing smoothness and time optimality. Among all the trajectories, the one with the least time to its goal is selected as the best trajectory and executed by the robot.
The following state (see Fig.3b and Fig.3d) follows a principle similar to that of the chasing state. The main difference is that the leader is now within the robot’s costmap. Instead of goal line sets in the chasing state, the candidate goal points are sampled on a circle centered at the leader, with a radius equal to the safe distance $D_{t}$. Similarly, these points are segmented by obstacles into several sub-goal circle sets $\mathcal{G}_{i}$.
During the optimization, the end point $p_{\tau}^{end}$ is constrained within the goal circle set $G_{k}$, and its orientation $\theta_{\tau}^{end}$ is constrained to head to the leader to reduce the risk of loss:
$\begin{gathered}\overrightarrow{\bar{p}_{l}\;\mathcal{G}_{k}^{start}}\times\overrightarrow{\bar{p}_{l}\;p_{\tau}^{end}}\geq 0,\hskip 20.0pt\overrightarrow{\bar{p}_{l}\;p_{\tau}^{end}}\times\overrightarrow{\bar{p}_{l}\;\mathcal{G}_{k}^{end}}\geq 0\\ |\overrightarrow{\bar{p}_{l}\;p_{\tau}^{end}}|=D_{t},\hskip 20.0pt|\theta_{\tau}^{end}-atan(\bar{p}_{l}-p_{\tau}^{end})|\leq\epsilon\end{gathered}$ (9)
@@
The planning state (see Fig.3c) is activated when the leader is no longer within the robot’s field of view. This typically occurs when the robot is detouring around an obstacle, or alternates with the following state when the leader is partially occluded. During the optimization, the endpoint of the trajectory $p_{\tau}^{end}$ is constrained to coincide with the last known position $\bar{p}_{l}$ and orientation $\bar{\theta}_{l}$ of the leader:
$\begin{gathered}|p_{\tau}^{end}-\bar{p}_{leader}|<=\epsilon,\hskip 20.0pt|\theta_{\tau}^{end}-\bar{\theta}_{leader}|<=\epsilon\end{gathered}$ (10)
@@
The retreating state (see Fig.3e) is activated when the leader actively approaches the robot, in which case the robot is expected to move backward to avoid potential collisions with the leader. The goal circle sets in this state are identical to those used in the following state. Among all points within the goal sets, the one closest to the robot is selected as the initial goal, and the endpoint of the trajectory is allowed to slide along the corresponding goal set $\mathcal{G}_{k}$. This sliding mechanism follows the same formulation as defined in Equ.9.
@@
The switching state (see Fig.3f) is activated when a new leader is assigned, most commonly triggered by a large language model [12]. In this state, both the memory buffer and the distance frame buffer are reset based on the current input. The system then directly transitions to the appropriate subsequent state according to the relative relationship between the robot and the newly assigned leader.
IV Simulation
The simulation evaluates the performance of the proposed leader-following and obstacle avoidance framework, Follow Everything, in a Gazebo simulation environment. The follower robot model is a legged robot, Unitree Go2, with a maximum speed of $V^{\text{max}}=1.5$ m/s and dimensions of 70x31x40 cm. The Realsense D435i is used for leader segmentation, and the Mid360 is employed to generate the point cloud. Four different leader scenarios are tested: a mobile robot as the leader in a playground environment, a UAV as the leader in a forest scenario, a pedestrian as the leader in a factory setting, and a stop sign as the leader in a dynamic scenario. In each test scenario, 10 leader trajectories are recorded using ROSbag and replayed to ensure a fair comparison. Each trajectory is tested four times to minimize the effects of environmental noise, resulting in 40 tests per scenario and 160 tests in total.
Four metrics [23] are considered: (1) the follow success rate (defined as successful when the robot maintains a safe distance from the leader and the leader remains in the robot’s FOV after the leader reaches its target), which is the most direct measure of the robot’s following robustness; (2) the average leader loss time ratio (the ratio of time during which the leader is lost from the robot’s view to the total test time), which reflects the causes of following failure in detail; (3) the collision rate (the ratio of collision occurrences to total 160 tests), which measures the following safety; and (4) the average distance (the average distance between the robot and the leader during the total 160 tests), where a larger distance can increase the risk of losing the leader.
Our method, Follow Everything, is compared against four baselines: (1) Alaa [13], which segments the leader using SAM2 and plans trajectories with a PID controller; (2) SA-MPC[21], which detects the leader using YOLO11 and utilizes an enhanced MPC-based planner; (3) FE-N-DFB, an ablated version of our framework that does not incorporate the distance frame buffer; and (4) FE-N-GP, another ablated version of our framework that replaces the graph-based planner with the MPC controller of SA-MPC.
Further details regarding the simulation setup and parameters are available in the source code [^1] https://follow-everything.github.io.
@@ IV Simulation
TABLE I: Comparison Result
@@
When the UAV flies directly over two closely spaced obstacles, as shown in Fig.5(a), the gap between them is too narrow for the follower to pass through. So our Follow Everything enables the robot to promptly detour around the obstacles. During this process, the UAV may temporarily disappear from the robot’s FOV. While the goal-state adaptation transitions from the following state to the planning state, the graph-based planner ensures behavioral consistency by switching the robot’s goal from a circle set to a point.
Figure 6: Simulation for following a pedestrian in a factory. (a–b) The robot follows the leader while navigating around obstacles. (c) Upon receiving a prompt specifying a new leader far away from the robot, the chasing state is activated, and the robot passes underneath the workstation to approach the new leader. (d) When the leader steps backward, the robot also retreats to maintain a safe distance. (e) When the leader turns a corner and leaves the robot’s FOV, the planning state is triggered to search for the leader.
After the UAV detours around two obstacles, as shown in Fig.5(b), the graph-based planner generates multiple feasible trajectories and selects the fastest one, rather than simply replicating the UAV’s historical path. This strategy helps the robot maintain a close distance to the leader. As shown in Table.LABEL:table and Fig.7, our method achieves the lowest average distance of 2.0 meters and robustly maintains it with the smallest variance, which reduces the likelihood of the leader leaving the robot’s FOV (too close can cause the UAV to leave the robot’s view, too far makes it hard to recognize), thus improving the overall follow success rate.
Figure 7: Average distance between the robot and the leader in 160 tests.
@@
Fig.6(d) illustrates the retreating state, where the leader actively steps backward. The robot is expected to retreat accordingly while avoiding nearby construction cones. This behavior significantly enhances safety. As a result, Follow Everything achieves a collision rate of only 1.8%, which is substantially lower than that of other methods. A key perception challenge in this state is that when the leader is close to the robot, only a portion of the leader’s body is visible. The YOLO-based method struggles to recognize such partial views as the leader, resulting in a high collision rate of 80.6%. Moreover, when the leader starts moving forward again, SAM2 without a distance frame buffer often fails to re-identify the leader, leading to a collision rate of 66.9% in Alaa. and 25.0% in FE-N-DFB.
Fig.6(e) shows how the planning state contributes to maintaining robust following behavior after the leader turns a corner and temporarily disappears from the robot’s FOV.
Figure 8: Illustration of real-world experiments in indoor (a-g) and outdoor (h-k) scenarios. (a-c) While following a pedestrian, maintaining an appropriate safe distance allows the robot to rotate in place and avoid unnecessary detours. (d-e) When following another robot, the leader is reliably identified and followed. (f-g) The planning state allows the robot to follow and re-identify the leader after moving out of the room. (h-i) Instead of replicating the leader’s historical trajectory, the robot takes a shorter path to quickly approach the leader. (j-k) The alternation between following and planning states enables the robot to follow the leader while avoiding obstacles when the leader steps over a box.
Stop sign leader in dynamic scenario. In this scenario, the leader is a stop sign, and dynamic obstacles move randomly, changing their velocities every 3 seconds. The graph-based planner enables the generation of multiple candidate trajectories, ensuring that a safer trajectory is not only available but also selected. This capability allows Follow Everything to significantly outperform other methods, achieving the lowest collision rate of 1.8%.
V Real-world Experiment
@@
Following state in indoor scenario. As shown in Fig.8(a), a suitcase obstacle is placed in the center of the scene, and the robot first follows the pedestrian, maintaining a relatively stable safe distance. As shown in Fig.8(b), during the pedestrian’s clockwise detour around the obstacle, the robot continuously stays in the following state, adjusting its orientation to always face the leader while maintaining a safe distance, thus avoiding unnecessary detours. As shown in Fig.8(c), once the pedestrian moves towards a further location, the robot detours counterclockwise around the obstacle and continues to follow the pedestrian. As shown in Fig.8(d-e), when the pedestrian leader is replaced by a legged robot with wheels, the robot’s following behavior remains stable.
@@
VI Conclusion
This paper presents a unified framework for robust and flexible leader-following. First, in the perception module, a segmentation-based model is employed instead of traditional detection methods, enabling the robot to follow any type or shape of leader. Furthermore, a distance frame buffer is integrated into the segmentation model to address the challenge of retracing the leader after temporary occlusions or short-term loss within the robot’s field of view. Second, a goal-aware adaptation is proposed, which dynamically adjusts parameters and generates trajectories appropriate goal states. This allows the robot to adapt its behavior according to the leader’s motion state, effectively reducing the risk of losing the leader or causing collisions.
Comprehensive simulations and real-world experiments were conducted using a legged robot as the follower and various objects as the leader—including UAVs, mobile robots, pedestrians, stop signs, and other legged robots—in both indoor and outdoor environments. The results demonstrate that the proposed approach greatly improves the success rate of following, reduces the duration of visual loss of the leader, reduces the collision rate, and decreases the average distance between the robot and the leader. Overall, the framework achieves higher levels of efficiency, safety, and robustness.

### v3

@@ II-D Goal-aware Adaptation
Chasing State (Fig. 3a) is activated when the leader is within the robot’s FOV but outside the costmap range, where the robot’s primary task is to navigate around nearby obstacles and chase the leader as quickly as possible. Unlike conventional methods that select the local goal as the intersection point between the global path and the costmap edge, the chasing state extends this point along the costmap edge to form a goal line of length $L_{t}$, which is proportional to the distance between the robot $\textbf{p}_{r}$ and the leader $\bar{\textbf{p}}_{l}$:
$L_{t}=\alpha(|\textbf{p}_{t}^{r}-\bar{\textbf{p}}_{t}^{l}|-\frac{W_{map}}{2}),$ (10)
@@
The whole goal line is separated at the locations where it intersects with obstacles, resulting in several sub-goal line sets $\mathcal{G}_{i}=\{\textbf{p}^{g}_{1},\textbf{p}^{g}_{2},...,\textbf{p}^{g}_{n_{i}}\}$, which are then added as additional nodes to the topological graph along with their corresponding connections. Therefore, the goal constraint $\mathcal{I}(\textbf{p}_{m}^{\tau},\mathcal{G}_{i})$ of Equ.(8) in this state is defined as:
$\begin{gathered}\overrightarrow{\textbf{p}^{g}_{1}\textbf{p}^{g}_{n_{i}}}\times\overrightarrow{\textbf{p}^{g}_{1},\textbf{p}^{\tau}_{m}}=0,\\ 0\leq\overrightarrow{\textbf{p}^{g}_{1}\textbf{p}^{g}_{n_{i}}}\cdot\overrightarrow{\textbf{p}^{g}_{1},\textbf{p}^{\tau}_{m}}\leq|\overrightarrow{\textbf{p}^{g}_{1}\textbf{p}^{g}_{n_{i}}}|^{2},\end{gathered}$ (11)
where $\textbf{p}^{g}_{1}$ and $\textbf{p}^{g}_{n_{i}}$ denote the endpoints of a goal set $\mathcal{G}_{i}$, and the $\textbf{p}^{\tau}_{m}$ is the last point on the trajectory $\tau$. The goal constraint in this state allow the point to $\textbf{p}^{\tau}_{m}$ slides along the goal set during optimization so that a faster trajectory can be optimized and obtained.
The velocity constraint $\mathcal{V}(\textbf{p}_{i}^{\tau},\textbf{p}_{i+1}^{\tau},V^{max})$ of Equ.(8) considers only the robot’s physical speed limit $V^{max}$ in this state. For each adjacent points, the constraint is defined as:
$||\textbf{p}_{i}^{\tau}-\textbf{p}_{i+1}^{\tau}||_{2}/\Delta T\leq V^{max},$ (12)
where $\Delta T$ is a fixed time interval. In this case, the robot can approach the goal as quickly as possible.
Following State (Fig. 3b and Fig. 3d) is activated when the leader is within the robot’s FOV and also within the current costmap. The primary task is to follow the leader while avoiding unnecessary detours, as in the experiment in Fig. 9b. A Kalman filter is used to estimate the leader’s state, where the Normalized Innovation Squared (NIS) [20] value serves as an indicator of prediction uncertainty. Based on this, a safe distance between the follower and the leader is dynamically adjusted as:
$D_{t}=\text{Clip}\left(\alpha\text{NIS},\,D^{min},\,D^{max}\right),$ (13)
where $D^{min}$ and $D^{max}$ are predefined lower and upper bounds, and $\alpha$ is a scaling factor. Instead of using goal line sets as in the chasing state, candidate goal points are sampled along a circle centered on the leader, with a radius equal to the safe distance $D_{t}$ in this state. These points are further segmented by obstacles into several sets of subgoal circles $\mathcal{G}_{i}$. In this state, the goal constraint $\mathcal{I}(\textbf{p}_{m}^{\tau},\theta_{m}^{\tau},\mathcal{G}_{i})$ at Equ.(8) allows the last trajectory point $\textbf{p}_{m}^{\tau}$ to slide along the goal circle set $\mathcal{G}_{i}=\{\textbf{p}^{g}_{1},\textbf{p}^{g}_{2},...,\textbf{p}^{g}_{n_{i}}\}$, while constraining the orientation $\theta_{m}^{\tau}$ to face the leader, thereby reducing the risk of losing visual contact:
@@
The maximum speed of the robot $V_{t}^{max}$ in this state is determined by the relative distance between the robot $\textbf{p}_{t}^{r}$ and the leader $\bar{\textbf{p}}_{t}^{l}$, as well as the estimated velocity of the leader $\bar{\textbf{v}}_{t}^{l}$. The velocity constraint $\mathcal{V}(\textbf{p}_{i}^{\tau},\textbf{p}_{i+1}^{\tau},V^{max})$ of Equ.(8) reduces the robot’s physical speed limit $V^{max}$ to $V^{max}_{t}$, avoiding frequent stop-and-go behavior of the robot. For each adjacent points, the constraint is defined as:
$\begin{gathered}V_{t}^{max}=\text{Clip}\left(\alpha_{1}|\bar{\textbf{v}}_{t}^{l}|+\alpha_{2}|\bar{\textbf{p}}_{t}^{l}-\textbf{p}_{t}^{r}|,\,0,\,V^{max}\right),\\ ||\textbf{p}_{i}^{\tau}-\textbf{p}_{i+1}^{\tau}||_{2}/\Delta T\leq V^{max}_{t},\end{gathered}$ (15)
where $\alpha_{1}$ and $\alpha_{2}$ are scaling factors. The function $\text{Clip}(\cdot)$ limits the value within the physical kinematics of the robot.
@@
Planning state (Fig. 3c) is activated when the leader is no longer within the robot’s field of view. This typically occurs when the robot is detouring around an obstacle or alternates with the following state when the leader is partially occluded. In this state, the goal constraint $\mathcal{I}(\textbf{p}_{m}^{\tau},\theta_{m}^{\tau},\bar{\textbf{p}}_{t}^{l})$ of Equ.(8) sets the trajectory endpoint $\textbf{p}_{m}^{\tau}$ and $\theta_{m}^{\tau}$ to be the last known position $\bar{\textbf{p}}_{t}^{l}$ and orientation $\bar{\theta}_{t}^{l}$ of the leader:
$\begin{gathered}|\textbf{p}^{\tau}_{m}-\bar{\textbf{p}}_{t}^{l}|<=\epsilon,\hskip 20.0pt|\theta_{m}^{\tau}-\bar{\theta}_{t}^{l}|<=\epsilon.\end{gathered}$ (16)
The velocity constraint $\mathcal{V}(\textbf{p}_{i}^{\tau},\textbf{p}_{i+1}^{\tau},V^{max})$ of Equ.(8) obeys the same way as that of chasing state, so that the robot can quickly approach the leader to reduce the risk of losing the visual contact.
@@
Retreating State (Fig. 3e) is activated when the leader actively approaches the robot. The robot is expected to move backward to avoid potential collisions with the leader. The velocity constraint $\mathcal{V}(\textbf{p}_{i}^{\tau},\textbf{p}_{i+1}^{\tau},V^{max}_{t})$ and goal constraints $\mathcal{I}(\textbf{p}_{m}^{\tau},\theta_{m}^{\tau},\mathcal{G}_{i})$ in this state are identical to those used in the following state. The one closest to the robot is selected as the initial goal, and the endpoint of the trajectory is allowed to slide along the corresponding goal set $\mathcal{G}_{i}$.
@@
Switching state (Fig. 3f) is activated when a new leader is assigned, most commonly triggered by a large language model [21]. In this state, both the memory buffer and the distance frame buffer are reset based on the current input. The system then directly transitions to the appropriate subsequent state according to the relative relationship between the robot and the newly assigned leader.
III Simulation
The simulation evaluates our Follow Everything in a Gazebo with a legged robot follower ($V^{max}=1.5,\mathrm{m/s}$), using a Realsense D435i for segmentation and a Mid360 for point clouds. Four scenarios are tested: a mobile robot as the leader in a playground, a UAV as the leader in a forest, a pedestrian as the leader in a factory, and a stop sign as the leader in a dynamic scenario. In each test scenario, $10$ leader trajectories are recorded using ROS bag and replayed to ensure a fair comparison. Each trajectory is tested four times to minimize the effects of environmental noise, resulting in $40$ tests per scenario and $160$ tests in total. Four metrics [13] are considered: (1) the follow success rate (the percentage of runs where the robot maintains a safe distance and keeps the leader in view after the leader reaches its target), (2) the average leader loss time ratio (the ratio of time the leader is out of view), (3) the collision rate (the ratio of collisions over 160 tests), and (4) the average distance (the average robot–leader distance, where larger values risk visual loss). Our method, Follow Everything, is compared against four baselines: (1) Alaa [10], which segments the leader using SAM2 and plans trajectories with a PID controller; (2) SA-MPC[11], which detects the leader using YOLO11 and utilizes an enhanced MPC-based planner; (3) FE-N-DFB, an ablation of our framework that does not incorporate the distance frame buffer; and (4) FE-N-GP, another ablation of our framework that replaces the graph-based planner with the MPC controller of SA-MPC. Further details regarding setups and parameters are available at the website.
Figure 5: Simulation for following a pedestrian in a factory. (a–b) The robot follows the leader when partially observed. (c) Upon receiving a prompt specifying a new leader, the chasing state is activated, and the robot passes underneath the workstation to approach the new leader. (d) When the leader steps backward, the robot retreats to maintain a safe distance. (e) When the leader turns a corner and leaves the robot’s FOV, the planning state is triggered.
@@ III Simulation
When the UAV flies over two closely spaced obstacles, as shown in Fig. 6(a), the gap between them is too narrow for the follower to pass through. So our Follow Everything enables the robot to promptly detour around the obstacles. During this process, the UAV may temporarily disappear from the robot’s FOV. While the goal-state adaptation transitions from the following state to the planning state, the graph-based planner ensures behavioral consistency by switching the robot’s goal from a circle set to a point.
TABLE I: Comparison of our Follow Everything to baselines and ablations. Our method achieves the best performance in all metrics
@@
After the UAV detours around two obstacles, as shown in Fig. 6(b), the graph-based planner generates multiple feasible trajectories and selects the fastest one, rather than replicating the UAV’s historical path. This strategy helps the robot maintain a close distance to the leader. As shown in Table. LABEL:table and Fig. 7, our method achieves the lowest average distance of 2.0 meters and robustly maintains it with the smallest variance, which reduces the likelihood of the leader leaving the robot’s FOV (too close can cause the UAV to leave the robot’s view, too far makes it hard to recognize), thus improving the overall follow success rate.
@@
Fig. 5(d) illustrates the retreating state, where the leader actively steps backward. The robot retreats accordingly while avoiding nearby construction cones, ensuring safety. As a result, Follow Everything achieves a collision rate of only 1.8%, which is substantially lower than that of other methods. A key perception challenge here is that when the leader is close to the robot, only a portion of the leader’s body is visible. The YOLO-based method struggles to recognize such partial views as the leader, resulting in a high collision rate of 80.6%. Moreover, when the leader starts moving forward again, SAM2 without a distance frame buffer often fails to re-identify the leader, leading to a collision rate of 66.9% in Alaa. and 25.0% in FE-N-DFB. Fig. 5(e) shows how the planning state contributes to maintaining robust following behavior after the leader turns a corner and temporarily disappears from the robot’s FOV.
Figure 7: Average robot-leader distance in 160 tests: Alaa 3.3 m, SA-MPC 2.4 m, FE-N-DFB 3.1m, FE-N-GP 2.6 m, and Follow Everything 2.0 m.
Stop sign leader in dynamic scenario. As shown in Fig.8, in this scenario, the leader is a stop sign, and dynamic obstacles move randomly, changing their velocities every 3 seconds. The graph-based planner enables the generation of multiple candidate trajectories, ensuring that a safer trajectory is not only available but also selected. This capability allows Follow Everything to significantly outperform other methods, achieving the lowest collision rate of 1.8%.
IV Real-world Experiment
@@
Following state in indoor scenario. As shown in Fig. 9(a), a suitcase is placed in the scene center, and the robot first follows the pedestrian, maintaining a relatively stable safe distance. As shown in Fig. 9(b), during the pedestrian’s clockwise detour around the obstacle, the robot continuously stays in the following state, adjusting its orientation to always face the leader while maintaining a safe distance, thus avoiding unnecessary detours. As shown in Fig. 9(c), once the pedestrian moves towards a further location, the robot detours counterclockwise around the obstacle and continues to follow the pedestrian. As shown in Fig. 9(d-e), when the pedestrian leader is replaced by a legged robot with wheels, the robot’s following behavior remains stable.
@@
V Conclusion and Future Work
This paper presents a unified framework for robust leader-following. A segmentation model with a distance frame buffer is introduced, enabling the robot to follow arbitrary leaders. A goal-aware adaptation and a graph-based planner are proposed, which dynamically adjust parameters, decide goal states, and generate trajectories. Simulations and experiments using a legged robot as the follower and UAVs, mobile robots, pedestrians, stop signs, and other legged robots as leaders in both indoor and outdoor environments demonstrate that the proposed approach greatly improves the success rate of following, reduces the duration of visual loss of the leader, reduces the collision rate, and decreases the average distance between the robot and the leader.
Figure 9: Illustration of real-world experiments. (a-c) While following a pedestrian, an appropriate safe distance allows the robot to rotate in place and avoid unnecessary detours. (d-e) The robot leader is reliably identified and followed. (f-g) The planning state allows the robot to follow and re-identify the leader after moving out of the room. (h-i) The robot takes a shorter path to quickly approach the leader. (j-k) The alternation between following and planning states enables the robot to follow the leader while avoiding obstacles when the leader steps over a box.